OpenStudy (automaticloveletter):

Solve by factoring. 4x^2-50x+49=50x

OpenStudy (gniiij):

ok so you want the equation to equal 0

OpenStudy (gniiij):

so get rid of the 50x

OpenStudy (automaticloveletter):

okay and

Directrix (directrix):

4x^2-50x+49=50x 4x² - 100x + 49 = 0

Directrix (directrix):

Try to factor this: 4x² - 100x + 49

Directrix (directrix):

(2x - ?) * (2x - ?) = 0